OpenStudy (anonymous):

3(x + 2) + 4(x - 5) = 10

OpenStudy (anonymous):

x = 3.428571429. k?

OpenStudy (uri):

3x+6+4x-20=10 7x-14-10=0 7x-24=0

OpenStudy (yanasidlinskiy):

Substitue the 3 into the parenthases. and the 4.